OpenStudy (anonymous):

For the function f(x)=(3-4x)^2, find the inverse. Determine whether the inverse is a function. I need to show my steps for this one as well @SolomonZelman

OpenStudy (solomonzelman):

okay, first step: replace the "f(x) with a y"

OpenStudy (solomonzelman):

do this for me please

OpenStudy (anonymous):

sorry i'm back

OpenStudy (anonymous):

y=(3-4x)^2

OpenStudy (solomonzelman):

it's alright. now, switch the y and x. (put y where x is, and put x where y is) you get... /

OpenStudy (anonymous):

x=(3-4y)^2

OpenStudy (solomonzelman):

yes. now take the square root of both sides (because we need to solve for y, but do this not by changing it back)

OpenStudy (anonymous):

\[\sqrt{x} =3-4\]

OpenStudy (solomonzelman):

you forgot something next to the 4, and the \(\large\color{black}{ \pm }\) next to the \(\large\color{black}{ \sqrt{x} }\)

OpenStudy (anonymous):

oh oops \[\pm \sqrt{x}=3-4x\]

OpenStudy (solomonzelman):

got the plus minus, but what is next to the 4?

OpenStudy (anonymous):

oh y not x

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ \pm \sqrt{x}=3-4y }\)

OpenStudy (solomonzelman):

subtract 3 from both sides.

OpenStudy (anonymous):

\[\pm \sqrt{x}-3=-4y\]

OpenStudy (solomonzelman):

now, divide both sides times -4.

OpenStudy (anonymous):

\[\frac{ \pm \sqrt{x}-3 }{ -4 }\]

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ y=(\pm \sqrt{x}-3)/(-4) }\)

OpenStudy (solomonzelman):

we can re-write it though.

OpenStudy (solomonzelman):

\(\large\color{black}{ y=(\pm \sqrt{x}-3)/(-4) }\) \(\large\color{black}{ y=(-1)(\pm \sqrt{x}-3)/(4) }\) \(\large\color{black}{ y=(\mp \sqrt{x}+3)/(4) }\) \(\large\color{black}{ y=(3\mp\sqrt{x})/(4) }\)

OpenStudy (solomonzelman):

and then, this would be same as, \(\large\color{black}{ y=(3\pm\sqrt{x})~/~4 }\)

OpenStudy (solomonzelman):

I can't scroll on it

OpenStudy (solomonzelman):

anyways, here is the graph. https://www.desmos.com/calculator/rdmctudroq

OpenStudy (solomonzelman):

is it a function?

OpenStudy (anonymous):

yes?

OpenStudy (solomonzelman):

does it pass the vertical line test?

OpenStudy (solomonzelman):

saying, is there (ever on the function) more than 1 point with the same x-coordinate, or on the same vertical line

OpenStudy (anonymous):

I dont know, when i graphth original it is goung up

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/yowqmozzja

OpenStudy (solomonzelman):

this is what I refer to a vertical line test. if I draw a vertical line, will 2 points on a function be on it?

OpenStudy (solomonzelman):

if 2 points (or more), that are on a function, are on a vertical line, then this function is NOT a function.

OpenStudy (solomonzelman):

if on any vertical line, there is only one point, then it is a function.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

I have to lave for about an hour.

OpenStudy (anonymous):

ok, is the screenshot right though with the steps before you go?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I couldn't scroll to see all steps on the screnshot

OpenStudy (solomonzelman):

I am near the comp, but I am doing some other activity now though...

OpenStudy (anonymous):

the only thing above is the initial equation nothing is needed to be scrolled. You can see everything i put in the thingy

OpenStudy (solomonzelman):

there are some mistakes, try to go over and repost.

OpenStudy (solomonzelman):

in your last step, you should have a positive 4 on the bottom